Soldatna Corporation's controller frequently prepared T-accounts to analyze inventory.  However, his penmanship was poor and he is no longer employed by the company.  The following was reconstructed from a scratch pad left on his desk.  The missing values (?) were illegible.  Analyze the information and answer the requirements that follow.
Raw Materials Cost of Goods Sold
beg. bal. ? ? 491,000 ?
175,000 491,000 ?
175,000 ?
481,000
102,000
Work in Process Factory Overhead
beg. bal. 44,000 460,000 35,000 ?
113,000 40,000
? 75,000
? ?
477,000 ? ? ?
? ?
Finished Goods
beg. bal. 131,000 ?
?
? ?
100,000
(a) Overhead is applied at 100% of direct labor cost.  How much was direct labor?
(b) Was overhead over- or underapplied, and by how much?
(c) Which inventory category increased?
(d) Other factory overhead, besides indirect material and indirect labor, was $75,000.  Indirect labor was 25% of the direct labor.  How much was indirect material?
(e) How much was cost of goods manufactured?
Answer #1

a) Cost of goods sold overapplied = 491,000 - 481,000 =10,000

Factory overhead = 35,000 + 40,000 + 75,000 + 10,000 = 160,000

Direct Labor = 160,000 *100% = 160,000

b) Overhead over applied = $10,000

c) Raw materials inventory category has increased.

d) Indirect labor = 160,000 - 75,000 - 10,000 - 160,000*25% = 35000

e) Cost of goods manufactured = 491,000 + 100,000 - 131,000 = 460,000